# Can LLM-based Financial Investing Strategies Outperform the Market in Long Run?

### v1

@@ 1 Introduction
Backtesting is the standard method for assessing investment strategies, simulating them on historical data to evaluate profitability and robustness (Chan, 2021). However, current LLM investing research suffers from fragmented, underdeveloped evaluation practices. Most studies assess performance over short periods, on few symbols, and often omit code release, limiting reproducibility. As summarised in Table 1, several recent methods evaluate over under a year, with fewer than ten stocks, and benchmark only against naïve baselines like Buy-and-Hold. Such short horizons and narrow stock universes lead to three well-documented sources of bias: survivorship bias (Garcia and Gould, 1993), where delisted or failed stocks are omitted; look-ahead bias (Chan, 2021), where future information inadvertently influences past decisions; and data-snooping bias (Bailey et al., 2015), where strategy performance is inflated through repeated testing on the same data. These biases can result in misleading performance assessments and undermine the validity of claimed improvements over traditional methods. This raises a central question: Can LLM-based investing strategies survive longer and broader robustness evaluations?
@@
    We propose FINSABER, the first comprehensive evaluation framework for LLM-based investing strategies that integrates 20 years of multi-source data, expands symbol coverage via unbiased selection, and mitigates survivorship, look-ahead, and data-snooping biases.
@@ 3 Definitions of Investing Strategies
3.1 Timing-based Investing Strategies
A timing-based investing strategy consists of predefined rules or algorithms that dictate buy, sell, or hold decisions for a given asset based on market data. The goal is to systematically anticipate short-term price movements to generate positive profit and loss (PnL) returns.
Let $P_{t}$ denote the adjusted closing price of a stock on day $t$, for $t=1,2,\dots,T$. A timing-based strategy determines a daily position $S_{t}\in\{-1,0,+1\}$, where $+1$ indicates a Buy, $-1$ a Sell, and 00 a Hold. Such strategies are typically driven by signals extracted from technical indicators, macroeconomic variables, sentiment, or machine learning models, aiming to exploit temporal inefficiencies in asset pricing.
3.2 Selection-based Investing Strategies
A selection-based strategy focuses on identifying a subset of assets from a broader universe to include in a portfolio. Unlike timing-based strategies, which focus on entry and exit points, selection-based approaches prioritise choosing the right assets based on their relative expected performance and the magnitude of the forecasts.
Rather than assigning precise portfolio weights, our setup considers equal-weighted or top-$k$ selection, where the aim is to allocate capital across high-conviction assets based on signals such as momentum, volatility, or model predictions. These strategies are periodically refreshed to adapt to changing market conditions.
The goal is to consistently favour assets expected to outperform, enabling long-term alpha through disciplined selection and rebalancing.
@@ 4 Why Broader and Longer Evaluation Matters
Also known as multiple testing bias, this occurs when repeated experimentation on the same dataset leads to overfitting. In finance, where sample sizes are small and the signal-to-noise ratio is very low, this bias is particularly problematic. Bailey et al. (2015) showed that evaluating many strategies on overlapping data inflates false positive rates, and that standard hold-out validation techniques often fail to guard against this issue.
@@
Symbol selection is another critical factor. Many LLM-based investing studies selectively use only a small number of well-known stocks such as TSLA and AMZN. These are both historical winners, which limits generalisability and embeds both survivorship and look-ahead bias into the evaluation. Omitting delisted or underperforming stocks distorts performance metrics and presents an incomplete picture of real-world investing conditions.
@@ 5 FINSABER
As discussed in §4, existing evaluations of LLM-based investors often suffer from survivorship bias, look-ahead bias, and data-snooping bias. These issues are largely due to limited evaluation periods and narrow stock selections. To address these limitations, we propose FINSABER—a backtesting framework for Financial INvesting Strategy Assessment with Bias mitigation, Expanded time, and Range of symbols.
FINSABER is composed of three core modules: a multi-source data module, a strategies base, and a bias-aware backtesting pipeline. Together, they support robust and reproducible benchmarking across four experimental setups described in §6. An overview is shown in Figure 1.
@@
FINSABER supports benchmarking across a wide range of investing strategies, including those driven by large language models. To do so, it integrates structured and unstructured financial data sources commonly used by LLM-based agents: historical stock prices, financial news, and company filings (10-K, 10-Q), all spanning from 2000 to 2024. To prevent look-ahead bias, all data inputs are aligned with each backtest window using only information available prior to the start date. Delisted stocks are included, and open-source equivalents are provided for reproducibility (more detail in Appendix A).
Modular Strategy Design and Bias-Mitigated Backtest Pipeline.
Strategies in FINSABER are categorised into selection-based (which pick stocks to trade) and timing-based (which decide when to trade) where timing-based strategies base includes rule-based, ML/DL, RL, and LLM investors and can be extended with custom modules (Appendix B). To mitigate survivorship bias, FINSABER uses historical S&P 500 constituent lists available at the beginning of each evaluation window and explicitly includes delisted stocks in the investable universe. To reduce data-snooping bias, the backtest pipeline employs rolling-window evaluation across multiple periods and asset selections. Both the window size and step are customisable, allowing researchers to simulate realistic evaluation settings across diverse market regimes. Timing-based strategies are paired with selection strategies to ensure broader symbol coverage and avoid overfitting to a narrow stock set or short horizon.
@@ 6 Experiments
Our experiments directly address the methodological flaws in current LLM-based investing evaluations identified in §4, particularly survivorship and data-snooping biases caused by selective stock choices and short evaluation periods. We show how these practices yield overly optimistic results and how FINSABER enables fairer assessment.
Specifically, our experiments are structured into two main parts: (1) Demonstrating pitfalls of selective evaluation: We replicate the originally reported performance results of recent LLM investors on selectively chosen periods and symbols. We then extend this evaluation period while keeping the original symbols unchanged, revealing how performance deteriorates over longer time horizons. (2) Establishing fair and robust comparisons: We leverage a composite approach, incorporating systematic stock-selection methods to explicitly mitigate survivorship and data-snooping biases, thus enabling fair and reliable assessments of LLM investor performance. We only consider go-long positions in trading, as existing LLM strategies do not model or support short-selling behaviour.
@@ 6.1 Pitfalls of Selective Evaluation
Revisiting Originally Reported Claims.
We begin by replicating prior setups from the literature, particularly evaluating the FinMem and FinAgent strategy on TSLA, NFLX, AMZN, and MSFT for the originally chosen period (6 October 2022 to 10 April 2023) by Yu et al. (2023). In addition to reproducing their results, we incorporate a broader set of benchmarks, including traditional rule-based, ML, and DL strategies. Since prior studies often omit critical details such as the exact risk-free rate and transaction costs, both of which are essential for realistic backtesting, we explicitly set a risk-free rate of 0.03 based on historical average rates, and adopt Moomoo’s [^2] https://www.moomoo.com/ca/support/topic10_122?lang=en-us standard US commission fee (\$0.0049/share, minimum \$0.99/order), which is widely comparable to other brokerage platforms, such as HSBC and TradeUp [^3] https://www.tradeup.com/pricing/detail.
@@
LLM FinMem (GPT-4o-mini) 0.593 10.520 -30.144 47.518
FinMem (GPT-4o) 0.587 10.710 -33.856 53.156
@@
LLM FinMem (GPT-4o-mini) -2.278 -2.460 -2.460 3.476
FinMem (GPT-4o) 1.543 41.167 -20.417 51.030
@@
LLM FinMem (GPT-4o-mini) 0.589 4.257 -2.746 9.875
FinMem (GPT-4o) -0.610 -13.764 -30.681 40.316
@@
LLM FinMem (GPT-4o-mini) -0.086 -1.679 -14.963 28.246
FinMem (GPT-4o) 0.307 -3.953 -45.462 89.114
@@
FinAgent 0.306 -8.605 -55.827 104.742
Table 2: Backtest performance of benchmark strategies over the selective period (2022-10-06 to 2023-04-10), as presented in Yu et al. (2023). “-” metrics indicate no trading activities were triggered. Top and second-best values are marked in red and blue.
Table 2 summarises these results. Our analysis reveals that LLM investors are not universally superior, even within their originally favoured setups. Specifically, FinMem shows significant outperformance only for TSLA, whereas traditional benchmarks remain competitive or superior for other symbols. These observations caution against overly optimistic interpretations from selective evaluations. Additionally, we find that LLM-based strategies exhibit high annual volatility and substantial maximum drawdowns, indicating a high-risk profile. This observation underscores the importance of explicit risk assessments in evaluating such strategies.
@@
To further illustrate the limitations of short evaluation horizons, we extend the evaluation period (2004–2024), while keeping the same four symbols (TSLA, NFLX, AMZN, MSFT). This longer historical perspective aims to assess the true robustness of the previously reported LLM outperformance.
@@
Table 3: Backtest performance of benchmark strategies for symbols from Yu et al. (2023) over extended period (2004-01-01 or earliest available to 2024-01-01). Top and second-best values are marked in red and blue.
Table 3 summarises these extended period results. Crucially, we observe that extending the evaluation horizon significantly diminishes the perceived superiority of LLM investors. Over two decades, traditional strategies like Buy and Hold consistently rank among the top performers across most symbols. For example, TSLA remains the only symbol where LLM investors (FinMem, FinAgent) retain clear superiority in annual return, whereas for NFLX, AMZN, and MSFT, Buy and Hold or other strategies outperform or match LLM strategies. This observation demonstrates that previously reported LLM advantages are likely short-lived, potentially hand-picked, and heavily dependent on the chosen evaluation period. Such fragility aligns with a central tenet in financial economics: asset prices are generally unpredictable in the long run, an assumption formalised in the Efficient Market Hypothesis (Fama, 1970). This foundational view cautions against over-interpreting transient outperformance as evidence of persistent edge, particularly when evaluation setups are narrow or selective.
@@ 6.2 Fair Comparisons with the Composite Approach
Table 4: Backtest performance under the Composite setup, using three different selection strategies across historical S&P 500 constituents (2004–2024), including delisted symbols. Top and second-best values are marked in red and blue.
@@
To overcome the aforementioned biases, we introduce the Composite evaluation setup within our FINSABER framework. This setup integrates systematic selection-based strategies to expand and diversify the stock universe, explicitly addressing survivorship and data-snooping biases. Specifically, we use three unbiased stock selection approaches (details in Appendix B.2): Random Five, Momentum Factor (Muller and and, 2010), and Volatility Effect (Blitz and Vliet, 2007).
@@
To effectively mitigate survivorship bias, our selection process uses only historical constituent lists of the S&P 500 available at the beginning of each evaluation period and explicitly includes delisted symbols. This ensures we avoid inadvertently selecting stocks that have survived due to their sustained performance over time. Moreover, the data-snooping bias is mitigated by evaluating a large and diversified universe of symbols (91, 84, and 63 symbols for Random Five, Momentum-based, and Volatility-based selection, respectively) over an extensive period of two decades, thus eliminating the biases associated with cherry-picking favourable stocks or short evaluation windows. These numbers refer to the total distinct symbols encountered across all rolling windows during evaluation, where we reselect stocks in each window.
@@
Table 4 summarises these comprehensive evaluations. Results obtained through this unbiased and systematic approach further validate our previous findings from the selected-four evaluation. Specifically, both the Random Five and Momentum-based selections reinforce the conclusion that the previously claimed superiority of LLM investors is largely driven by selective evaluation setups. For instance, in the Random Five setup, Buy and Hold, ATR Band and ARIMA outperform FinMem and FinAgent in terms of risk-adjusted metrics. Similarly, ARIMA and simple rule-based strategies often perform better than LLM-based methods under the Momentum-based selection. In the Volatility-based selection, traditional methods dominate even more clearly: Buy and Hold achieves the highest Sharpe (0.703), stability (1.291), and annual return (7.898%), while PPO and ARIMA again show strong all-round performance. LLM-based methods lag behind, with FinAgent offering moderate returns but lower Sharpe (0.241) and larger drawdowns. Notably, our reported LLM performances do not adjust for potential data leakage: given the use of pretrained models like GPT-4o, the LLMs may have seen parts of the data during training—yet they still fail to outperform traditional strategies under fair evaluation, casting further doubt on their real-world advantage.
@@
Nevertheless, it is important to acknowledge that LLM-based strategies still show potential regarding absolute annual returns. For instance, FinAgent achieves the highest annual return (13.950%) in the Momentum-based selection setup. However, the relatively weaker performance observed in Sharpe ratio (0.104) and maximum drawdown metrics suggests a clear need for improved risk management within LLM-driven approaches before they can be reliably adopted in practice.
@@ 7 Market Regime Analysis
A remaining question in evaluating LLM-based investing strategies is understanding whether strategies adapt appropriately across varying market conditions. As Kim et al. (2011) demonstrated, financial markets exhibit time-varying predictability and uncertainty across different economic, financial, and political regimes. Some strategies may be able to exploit these variations, while others may struggle to adapt effectively. Different market environments including bull, bear, and sideways, pose distinct challenges and opportunities: bull markets reward aggressive positioning and substantial exposure, bear markets require risk management, and sideways markets test strategies’ ability to navigate uncertainty without clear trends. By decomposing performance into these regimes, we can identify whether strategies are overly conservative, missing opportunities during bullish periods, or excessively aggressive, suffering large losses during downturns. Clarifying these regime-specific behaviours is crucial for interpreting the strengths and weaknesses of LLM-based investing strategies and for informing their future development (Hui and Chan, 2018).
@@
To analyse regime-specific performance, we use our composite setup under three selection strategies: Random 5, Momentum Factor (Muller and and, 2010), and Volatility Effect (Blitz and Vliet, 2007). For each timing strategy, we retrieve the Sharpe ratio within each 1-year window from Table 4. These are then averaged per {strategy, regime} pair to produce stable performance indicators across market conditions. Figure 2 illustrates the results, with green indicating strong Sharpe ratio and red signifying the opposite.
Traditional rule-based and predictor-based methods still set the standard. ATR Band, Turn of the Month and ARIMA deliver positive Sharpe in every regime, while Buy and Hold, our passive yardstick, posts 0.61 in bulls, 0.48 in sideways markets and only -0.28 in bears. No active strategy surpasses this passive Sharpe ratio in the bull regime, suggesting that many strategies, including the LLM ones, may struggle to fully capitalise on strong up-trends.

### v2

@@ 1 Introduction
Backtesting is the standard method for assessing investment strategies, simulating them on historical data to evaluate profitability and robustness (Chan, 2021). However, current LLM investing research suffers from fragmented, underdeveloped evaluation practices. Most studies assess performance over short periods, on few stock symbols, and often omit code release, limiting reproducibility. As summarised in Table 1, several recent methods evaluate over under a year, with fewer than ten stocks, and benchmark only against naïve baselines like Buy-and-Hold. Such short horizons and narrow stock universes lead to three well-documented sources of bias: survivorship bias (Garcia and Gould, 1993), where delisted or failed stocks are omitted; look-ahead bias (Chan, 2021), where future information inadvertently influences past decisions; and data-snooping bias (Bailey et al., 2015), where strategy performance is inflated through repeated testing on the same data. These biases can result in misleading performance assessments and undermine the validity of claimed improvements over traditional methods. This raises a central question: Can LLM-based investing strategies survive longer and broader robustness evaluations?
@@
    We propose FINSABER, the first comprehensive evaluation framework for LLM-based investing strategies that supports 20 years of multi-source data, including unstructured inputs such as news and filings, expands symbol coverage via unbiased selection, and mitigates survivorship, look-ahead, and data-snooping biases.
@@ 3 Definitions of Investing Strategies
Timing-Based Strategies
Timing-based strategies generate daily Buy ($+1$), Sell ($-1$), or Hold (00) signals based on market data such as prices, technical indicators, or model outputs. The objective is to capture short-term price movements through systematic trading rules.
Selection-Based Strategies
Selection-based strategies identify subsets of assets expected to outperform based on ranking signals. Assets are selected periodically using top-$k$ or thresholding. These strategies focus on cross-sectional alpha rather than timing individual trades.
@@ 4 Why Broader and Longer Evaluation Matters
Also known as multiple testing bias, this occurs when repeated experimentation on the same dataset leads to overfitting. In finance, where sample sizes are small and the signal-to-noise ratio is very low, this bias is particularly problematic. Bailey et al. (2015) showed that evaluating strategies on overlapping data inflates false positive rates, and that standard hold-out validation techniques often fail to guard against this issue.
@@
Stock selection is another critical factor. Many LLM-based investing studies selectively use only a small number of well-known stocks such as TSLA and AMZN. These are both historical winners, which limits generalisability and embeds both survivorship and look-ahead bias into the evaluation. Omitting delisted or underperforming stocks distorts performance metrics and presents an incomplete picture of real-world investing conditions.
@@ 5 FINSABER
As discussed in §4, existing evaluations of LLM-based investors suffer from survivorship bias, look-ahead bias, and data-snooping bias. These issues are largely due to limited evaluation periods and narrow stock selections. In this study, all subsequent findings and analyses are derived from our meticulously constructed backtesting framework, FINSABER [^2] Financial INvesting Strategy Assessment with Bias mitigation, Expanded time, and Range of symbols, which systematically addresses biases and is specifically designed to meet the practical needs of LLM-based strategies, including the integration of unstructured, multi-source data.
FINSABER comprises three core modules: (1) a multi-source data module, (2) a modular strategy base, and (3) a bias-aware two-step backtesting pipeline. Figure 1 illustrates the framework.
@@
LLM-based investing strategies utilise both structured and unstructured data such as historical stock prices, financial news, and company filings (10-K, 10-Q), spanning from 2000 to 2024. To prevent look-ahead bias, all data inputs are aligned with each backtest window using only information available prior to the start date. Survivorship bias is addressed by explicitly including delisted stocks, and open-source equivalents are provided for reproducibility (more detail in Appendix A).
Two-Step Pipeline for Bias Mitigation.
FINSABER applies a two-step pipeline. First, selection-based strategies operate on regularly updated, historically accurate constituent lists, for example, the S&P 500 including delisted symbols, at each window. This further mitigates survivorship bias from the stock selection process, ensuring the evaluation is not restricted to a limited or selectively surviving set of stocks. Subsequently, timing-based strategies which covers rule-based, ML, RL, and LLM-driven approaches will be used to execute daily trading decisions. The modular strategy base is easily extensible for custom methods (see Appendix B). To mitigate data-snooping bias, rolling-window evaluations are performed over diverse and dynamically changing asset selections and extended time horizons. Window size and step are customisable, enabling realistic simulation across different market regimes. Together, this pipeline ensures broad symbol coverage and prevents overfitting to narrow datasets or short evaluation horizons.
@@ 6 Experiments
Our experiments address methodological flaws in prior LLM-based investing evaluations identified in §4, specifically survivorship and data-snooping biases from selective stock choices and short evaluation periods. We demonstrate how these practices inflate results and illustrate how FINSABER enables fairer assessments.
Specifically, our experiments include two parts: (1) Pitfalls of selective evaluation: Replicating previously reported results on select periods and symbols, then extending this evaluation period to demonstrate performance deterioration. (2) Fair and robust comparisons: Implementing systematic stock-selection methods to explicitly mitigate survivorship and data-snooping biases for fairer LLM assessments. We only consider go-long positions, aligning with current LLM strategies.
@@ 6.1 Pitfalls of Selective Evaluation
Revisiting Reported Claims.
We begin by replicating earlier evaluation setups that demonstrated the effectiveness of LLM investing strategies on TSLA, NFLX, AMZN, and MSFT during the previously reported period (6 October 2022 to 10 April 2023). Additionally, we incorporate broader benchmarks, including traditional rule-based, ML, and DL methods. Previous studies omit key details such as exact risk-free rates and transaction costs. Thus, we explicitly set a historical average risk-free rate of 0.03 and use Moomoo’s [^3] https://www.moomoo.com/ca/support/topic10_122?lang=en-us standard US commission fee (\$0.0049/share, minimum \$0.99/order), comparable to HSBC and TradeUp [^4] https://www.tradeup.com/pricing/detail platforms.
@@
LLM FinMem (GPT-4o-mini) 0.927 19.940 -30.144 48.638
FinMem (GPT-4o) 0.404 5.312 -36.351 54.434
@@
LLM FinMem (GPT-4o-mini) 1.704 32.549 -13.018 34.766
FinMem (GPT-4o) 0.896 16.244 -15.234 38.209
@@
LLM FinMem (GPT-4o-mini) 0.297 2.800 -2.744 10.247
FinMem (GPT-4o) -0.968 -20.091 -31.164 40.896
@@
LLM FinMem (GPT-4o-mini) -0.554 -7.104 -14.588 25.969
FinMem (GPT-4o) 0.792 12.834 -13.555 33.884
@@
FinAgent 0.313 -10.598 -56.641 111.434
Table 2: Backtest performance of benchmark strategies over the previously reported period (2022-10-06 to 2023-04-10) where LLM investing strategies were shown to be effective. “-” metrics indicate no trading activities were triggered. Top in red and second-best in blue.
Table 2 summarises these results. Our analysis reveals that LLM investors are not universally superior, even within their originally favoured setups. Specifically, FinMem shows significant outperformance only for TSLA, whereas traditional benchmarks remain competitive or superior for other symbols. These observations caution against overly optimistic interpretations from selective evaluations. FinAgent, another LLM-based method, performs comparably to FinMem on NFLX and MSFT but generally does not offer consistent improvements across the set. Additionally, we find that LLM-based strategies exhibit high annual volatility and substantial maximum drawdowns, indicating a high-risk profile. This observation underscores the importance of explicit risk assessments in evaluating such strategies.
@@
To further illustrate the limitations of short evaluation horizons, We extend the evaluation period (2004–2024) using the same four symbols (TSLA, NFLX, AMZN, MSFT) to assess LLM performance robustness over the long term.
@@
Table 3: Backtest performance of benchmark strategies for previously reported LLM-selected symbols over an extended period (2004-01-01 or earliest available to 2024-01-01). Top in red and second-best in blue.
Table 3 summarises these extended period results. Crucially, extending the evaluation horizon significantly diminishes the perceived superiority of LLM investors. Over two decades, traditional strategies like Buy and Hold consistently rank among the top performers across most symbols. TSLA is the only case where LLM investors (FinMem, FinAgent) clearly lead in AR, while for NFLX, AMZN, and MSFT, Buy and Hold or other strategies match or outperform them. This suggests previously reported LLM advantages are likely short-lived, potentially hand-picked, and highly sensitive to the evaluation period.
@@ 6.2 Fair Comparisons with the Composite Approach
Table 4: Backtest performance under the Composite setup, using three different selection strategies across historical S&P 500 constituents (2004–2024), including delisted symbols. Top in red and second-best in blue.
@@
To overcome the aforementioned biases, we introduce the Composite evaluation setup within FINSABER. This setup integrates systematic selection-based strategies to expand and diversify the stock universe, explicitly addressing survivorship and data-snooping biases. Specifically, we use three unbiased stock selection approaches (details in Appendix B.2): Random Five, Momentum Factor (Muller and and, 2010), and Volatility Effect (Blitz and Vliet, 2007). Each selection strategy is executed at the start of every rolling window, serving as a rebalancing mechanism.
@@
To mitigate survivorship bias, we use historical constituent lists, specifically S&P 500 for US market, at each evaluation period’s start and explicitly include delisted symbols. To address data-snooping bias, we evaluate a large and diversified symbol universe: 91, 84, and 63 total distinct symbols for Random Five, Momentum-based, and Volatility-based selection, respectively. These counts reflect all unique symbols encountered across rolling windows, where stocks are reselected in each window, preventing cherry-picking and short-horizon bias.
@@
Table 4 summarises these comprehensive evaluations. Results obtained through this unbiased and systematic approach further validate our previous findings from the selected-four evaluation. Specifically, both the Random Five and Momentum-based selections reinforce the conclusion that the previously claimed superiority of LLM investors is largely driven by selective evaluation setups. For instance, in the Random Five setup, Buy and Hold, ATR Band and ARIMA outperform FinMem and FinAgent in terms of risk-adjusted metrics. Similarly, ARIMA and simple rule-based strategies often perform better than LLM-based methods under the Momentum-based selection. In the Volatility-based selection, traditional methods dominate even more clearly: Buy and Hold achieves the highest Sharpe (0.703), stability (1.291), and AR (7.898%), while PPO and ARIMA again show strong all-round performance. LLM-based methods lag behind, with FinAgent offering moderate returns but lower Sharpe (0.241) and larger drawdowns. Notably, our reported LLM performances do not adjust for potential data leakage: given the use of pretrained models like GPT-4o, the LLMs may have seen parts of the data during training—yet they still fail to outperform traditional strategies under fair evaluation, casting further doubt on their real-world advantage.
@@
Nevertheless, it is important to acknowledge LLM-based strategies still show potential regarding absolute annual returns. For instance, FinAgent achieves the highest AR (13.950%) in the Momentum-based selection setup. However, the relatively weaker performance observed in SPR (0.104) and MMD metrics suggests a clear need for improved risk management within LLM-driven approaches before they can be reliably adopted in practice.
@@ 7 Market Regime Analysis
Another key question in evaluating LLM-based investing strategies is whether they adapt appropriately across varying market conditions. Financial markets exhibit time-varying predictability and uncertainty across different economic, financial, and political regimes (Kim et al., 2011). Some strategies may exploit these variations, while others may struggle to adapt. Distinct market environments—bull, bear, and sideways—present unique challenges and opportunities: bull markets reward aggressive positioning and high exposure, bear markets require effective risk management, and sideways markets test a strategy’s ability to navigate uncertainty in the absence of clear trends. By decomposing performance across these regimes, it is possible to determine whether strategies are overly conservative and miss opportunities during bullish periods, or excessively aggressive and incur significant losses during downturns. Understanding these regime-specific behaviours is essential for interpreting the strengths and weaknesses of LLM-based investing strategies and for guiding their future development (Hui and Chan, 2018).
@@
To analyse regime-specific performance, we employ our composite setup using the three selection strategies outlined in §6.2. For each timing strategy, we retrieve the SPR within each 1-year window from Table 4. These are then averaged per {strategy, regime} pair to produce stable performance indicators across market conditions. Figure 2 illustrates the results, with green indicating strong SPR and red signifying the opposite.
Traditional rule-based and predictor-based methods still set the standard. ATR Band, Turn of the Month and ARIMA deliver positive Sharpe in every regime, while Buy and Hold, our passive yardstick, posts 0.61 in bulls, 0.48 in sideways markets and only -0.28 in bears. No active strategy surpasses this passive SPR in the bull regime, suggesting that many strategies, including the LLM ones, may struggle to fully capitalise on strong up-trends.
@@
8 Findings and Takeaways
The evaluations in §6 and §7 reveals structural limitations in current LLM-based investing strategies. These findings expose methodological flaws in recent studies and highlight the need for more rigorous and domain-aware evaluation.
First, the fragility of LLM-derived alpha under realistic conditions aligns with the Efficient Market Hypothesis (EMH) (Fama, 1970): asset prices reflect all available information, making it difficult to consistently achieve above-average returns through information-based trading. Despite access to large textual corpora, current LLMs fail to extract signals that persist beyond short windows, suggesting that observed gains may stem from narrow evaluation setups rather than genuine alpha.
Second, scaling laws observed in NLP and other domains do not translate effectively to financial applications. While larger models often yield better performance in NLP tasks (Kaplan et al., 2020), our results indicate that increasing model size does not reliably improve investing performance. For instance, FinMem with GPT-4o does not outperform its smaller variant. This suggests that financial markets impose intrinsic limits on extractable signals, regardless of model capacity (Harvey and Liu, 2013). Moreover, simple models like ARIMA or rule-based systems often outperform LLMs on risk-adjusted metrics. This indicates that complexity without financial logic such as trend detection, volatility targeting, or drawdown control, adds little value (and, 2001; Kabiri et al., 2023).
Third, LLMs poorly adapt to market regimes. They tend to be overly conservative in bull markets and excessively aggressive in downturns, contradicting the Adaptive Markets Hypothesis (AMH), which stresses dynamic risk management across shifting conditions (Lo, 2004).
Finally, biased evaluation designs, such as short evaluation periods and selectively chosen symbols, systematically exaggerate results. In finance, this is not merely a technical flaw but a practical risk, as misleading models can drive poor allocation decisions (Sullivan et al., 1999). FINSABER addresses this by enforcing rigorous, bias-aware benchmarking, setting a baseline for credible evaluation in the financial LLM field.